Complete the ordered pairs so that each is a solution for the given equation.
3x+4y=12 (0,_), (_,3/4), (_,0), (8/3,_) I do not have a clue as to where to start with this problem. I would really appreciate it if someone would please explain it step by step to help me understand the problem.

Answer by tutorcecilia(2144)

In ordered pairs, the first number is always the x-term and the second number is always the y-term. So (x, y) can be (1, 0) and the x=1; y=0
.
In this problem, identify either the x or the y-term, plug it into the problem and solve for the missing term:
.
(0,_), so x=0 and the y-term is missing:
3x+4y=12
3(0)+4y=12
0+4y=12
4y=12
4y/4=12/4
y=3
So, the point is (0,3)
.
(_,3/4), so the x-term is missing and y=3/4
3x+4(3/4)=12
3x+3=12
3x+3-3=12-2
3x=9
3x/3=9/3
x=3
So, the point is (3, 3/4)